Grade 12Physical Chemistry

sample of MnSO​4.4H2O is strongly heated in air. The residue (Mn3O4) left was dissolved in 100 ml of 0.1 N FeSO4containing dil. H2SO4. This solution was completely reacted with 50 ml of KMnO4Solution : 25 ml of this KMnO4 solution was completely reduced by 30 ml of 0.1 N FeSO4 Solution: Calculate the amountr of MnSO4.4H2O in sample.

RENU MISHRA
7 Years agoGrade 12

1 Answer

Arun
6 Years ago
MnSO 
4
​ 
 ⋅4H 
2
​ 
 O 
Δ
​ 
 Mn 
3
​ 
 O 
4
​ 
 
The residue Mn 
3
​ 
 O 
4
​ 
  is dissolved in FeSO 
4
​ 
 , which is reduced from Mn 
3
8
​ 
 +
  to Mn 
2+
 ,
 i.e., (Mn 
3
8
​ 
 +
 ) 
3
​ 
 +2e⟶3Mn 
2+
   
​ 
 ∵E 
Mn 
2+
 
​ 
 = 
3
2
​ 
 
M
​ 
  
​ 
 
For the reaction of KMnO 
4
​ 
  with FeSO 
4
​ 
 ,
Meq.ofKMnO 
4
​ 
 =Meq.ofFeSO 
4
​ 
 
25×N=30×0.1
∴N 
KMnO 
4
​ 
 
​ 
 = 
25
3
​ 
 
Meq. of FeSO 
4
​ 
  added to Mn 
3
​ 
 O 
4
​ 
 =100×0.1=10
Meq. of FeSO 
4
​ 
  left after reaction with Mn 
3
​ 
 O 
4
​ 
 =Meq.ofKMnO 
4
​ 
 used=50× 
25
3
​ 
 =6
∴ Meq. of FeSO 
4
​ 
  used for Mn 
3
​ 
 O 
4
​ 
 =10−6=4
∴ Meq. of MnSO 
4
​ 
 = Meq. of Mn 
3
​ 
 O 
4
​ 
 =4
∵ As reported above E 
MnSO 
4
​ 
 ⋅4H 
2
​ 
 O
​ 
 = 
2
3M
​ 
 = 
2
223×3
​ 
 
∴ 
2
3M
​ 
 
MnSO 
4
​ 
 
​ 
 
​ 
 ×1000=4
∴w 
MnSO 
4
​ 
 
​ 
 = 
2×1000
4×3×223
​ 
 =1.338g